OpenStudy (anonymous):

Factor the rational expression: 3t^4-9t^3/6t^2 thanks for the help! please show me how! i do not get it

OpenStudy (anonymous):

this is hard let me think

OpenStudy (anonymous):

ok

OpenStudy (mertsj):

\[\frac{3t^4-9t^3}{6t^2}=\frac{3t^4}{6t^2}-\frac{9t^3}{6t^2}\]

OpenStudy (mertsj):

Think of it like that and reduce each fraction.

OpenStudy (anonymous):

how about simplifying it? would it be the same answer?

OpenStudy (anonymous):

and thanks for your help so far

OpenStudy (mertsj):

yes

OpenStudy (mertsj):

yw

OpenStudy (anonymous):

ok. thanks!!